OpenStudy (anonymous):

You are planning to draw 2 cards at random without replacement. Your goal is to pull out a face card on the first draw and a heart in the second draw.

OpenStudy (istim):

(52P4)*(52P13) I am not sure thought.

OpenStudy (istim):

Do you know if that is how "Permutations" work?

OpenStudy (anonymous):

im not sure...

OpenStudy (anonymous):

i think i am supposed to use the multiplication rule

OpenStudy (istim):

I do not remember then.

OpenStudy (anonymous):

If the occurence of Event A changes the probability of Event B, then Events A and B are dependent. On the other hand, if the occurence of Event A does not change the probability of Event B, then Events A and B are independent.

OpenStudy (istim):

Well, the even does affect the second event. With one card drawn, the total number of cards in the deck decrease from 52 to 51.

OpenStudy (anonymous):

yeah that is right

OpenStudy (istim):

Yayy! Dumbcow is with us.

OpenStudy (istim):

52P4 Means that we took out 4 cards in order from 52, so that can't be right.

OpenStudy (dumbcow):

are we trying to find probability of drawing a face card 1st and a heart 2nd ?

OpenStudy (anonymous):

yes

OpenStudy (istim):

I think so.

OpenStudy (anonymous):

i really don't know how to do this one...

OpenStudy (dumbcow):

there are 2 cases then: case 1 -> the face card is a heart there are only 3 cards that fit (jack hearts, queen hearts, king hearts) then that leaves only 12 hearts remaining for next card \[P = \frac{3}{52}*\frac{12}{51}\] case 2-> face card is not a heart \[P = \frac{9}{52}*\frac{13}{51}\] to find combined probability, add the probabilities from both cases \[\rightarrow \frac{3*12}{52*51}+\frac{9*13}{52*51} = \frac{9}{15,028}\]

OpenStudy (dumbcow):

oh my final answer is off..i did some bad computation there sorry

OpenStudy (anonymous):

that makes sense

OpenStudy (dumbcow):

final answer: \[\rightarrow \frac{3}{52}\]

OpenStudy (anonymous):

great how about the quarter section...

OpenStudy (dumbcow):

well there is only 1 special quarter in the roll so probability of selecting it is 1/n where n is number of quarters \[\rightarrow \frac{1}{n} = \frac{3}{52}\]

OpenStudy (anonymous):

thanks you have been such a great help!!!!

OpenStudy (dumbcow):

your welcome it doesn't come out to a whole number though so i may have done something wrong, not sure

OpenStudy (anonymous):

thanks any ways...